OpenStudy (anonymous):

-2x^2-x+3 <--- find x-intercept by factoring and zero product property...thank you

OpenStudy (zpupster):

factor the equation first: (-2x -3)( x -1) then set each to 0 -2x-3) = 0 x-1=0 x=-3/2 and x=1

OpenStudy (anonymous):

and does that include zero product property?

OpenStudy (zpupster):

setting the factors to 0 is the zero product property

OpenStudy (anonymous):

gotcha thanks mate